# SCA: Highly Efficient Semantic-Consistent Unrestricted Adversarial Attack

### v5

@@ 5 Experiments
SAE[19] 0.7947 0.9492 15.6864 0.3767
ADef[1] 0.7856 0.9112 14.7823 0.3996
cAdv[2] 0.8398 0.9924 19.9874 0.2695
tAdv[2] 0.8366 0.9876 19.0012 0.2377
ACE[72] 0.8318 0.9891 19.7205 0.1799
ColorFool[47] 0.8448 0.9358 9.7676 0.5102
NCF[66] 0.8848 0.8135 14.3670 0.3194
AdvST[57] 0.8752 0.8144 14.2011 0.2908
ACA[10] 0.7442 0.5137 11.0014 0.5971
@@ 5.1 Experimental Setup
Datasets. We carry out our experiments using the ImageNet-compatible dataset [25], which includes 1,000 images from the ImageNet validation set [12]. This dataset has been frequently employed in related work [15, 16, 62, 66, 10].
Semantic Consistency Evaluation. We use CLIP Score [44], SSIM [58], PSNR [23], and LPIPS [67] to measure the semantic consistency between the generated unrestricted adversarial examples and the clean image from multiple aspects.
Attack Evaluation. We compare SCA with other unrestricted adversarial attacks, including SAE [19], ADef [1], cAdv [2], tAdv [2], ACE [72], ColorFool [47], NCF [66], AdvST [57], and ACA [10]. The default parameter settings for these attacks are used. The evaluation metric employed is attack success rate (ASR), which measures the proportion of images misclassified by the target models.
@@
Implementation Details. All experiments are performed using PyTorch on an NVIDIA Tesla A100. The DDPM steps are set to $T=10\to 20$, with attack iterations $N_{a}=10$, $\eta=0.04$, $\kappa=0.1$, and $\mu=1$. The version of Stable Diffusion [45] used is v1.5, and image captions are generated automatically via LLaVA-NeXT.
@@ 5.2 Semantic Consistency Comparison
5.3 Comparison of Attack Success Rates
We systematically evaluate the adversarial robustness of convolutional neural networks (CNNs) and vision transformers (ViTs) using several prominent adversarial methods: SAE [19], ADef [1], cAdv and tAdv [2], ACE [72], ColorFool [47], NCF [66], AdvST [57], ACA [10], and our proposed SCA. These methods were applied to MobileNetV2 (MN-v2), ResNet-50 (RN-50), MobViT-s, and ViT-B, with performance measured using the Average Attack Success Rate (%), which captures the transferability of adversarial examples across non-substitute models.
0ptclean imageRefer to caption 0ptcAdvRefer to caption 0ptSAERefer to caption 0ptADefRefer to caption 0ptColorFoolRefer to caption 0ptNCFRefer to caption 0ptAdvSTRefer to caption 0ptACARefer to caption 0ptOurs
@@ 5.3 Comparison of Attack Success Rates
0ptclean imageRefer to caption 0ptcAdvRefer to caption 0ptSAERefer to caption 0ptADefRefer to caption 0ptColorFoolRefer to caption 0ptNCFRefer to caption 0ptAdvSTRefer to caption 0ptACARefer to caption 0ptOurs
@@
5.4 Image Quality Comparison.
In this section, we conduct a detailed quantitative assessment of the image quality for adversarial examples, following the approach taken in prior works such as [47], [66], and [10]. Specifically, we employ non-reference perceptual image quality metrics to evaluate the visual realism and aesthetic appeal of the adversarial images generated by different methods. The metrics we use include NIMA [52], HyperIQA [50], MUSIQ [24], and TReS [17], which are widely recognized for their ability to assess image quality in the absence of a ground truth reference.
To provide a robust evaluation, we follow the PyIQA framework and apply models trained on large-scale aesthetic datasets. NIMA-AVA and MUSIQ-AVA are trained on the AVA [40] dataset, while MUSIQ-KonIQ is trained on the KonIQ-10K [20] dataset, which ensures a comprehensive evaluation of the generated adversarial examples. As shown in Table 3, our method consistently achieves image quality scores that are comparable to ACA, outperforming other methods such as ColorFool and NCF. This superior performance is likely attributable to the fact that both SCA and ACA leverage Stable Diffusion, a state-of-the-art generative model known for its ability to produce highly realistic and visually appealing images.
@@
Additionally, it is worth noting that the no-reference image quality metrics we employ are often trained on aesthetic datasets, many of which contain post-processed images, including those edited with tools such as Photoshop. These post-processed images are more aligned with human aesthetic preferences, which is an important consideration for assessing adversarial examples. As [10] points out, the similarity of our method to post-processing and image-editing techniques helps explain why our adversarial examples receive higher quality scores from these metrics. Our approach effectively functions as a form of image editing, subtly modifying the image in ways that retain or even enhance its aesthetic appeal, leading to better evaluation results.

### v6

@@ 5 Experiments
SAEhosseini2018semantic 0.7947 0.9492 15.6864 0.3767
ADefalaifari2018adef 0.7856 0.9112 14.7823 0.3996
cAdvbhattad2019unrestricted 0.8398 0.9924 19.9874 0.2695
tAdvbhattad2019unrestricted 0.8366 0.9876 19.0012 0.2377
ACEzhao2020adversarial 0.8318 0.9891 19.7205 0.1799
ColorFoolshamsabadi2020colorfool 0.8448 0.9358 9.7676 0.5102
NCFyuan2022natural 0.8848 0.8135 14.3670 0.3194
AdvST10292904 0.8752 0.8144 14.2011 0.2908
ACAchen2024content 0.7442 0.5137 11.0014 0.5971
@@ 5.1 Experimental Setup
Datasets. We carry out our experiments using the ImageNet-compatible dataset kurakin2018adversarial , which includes 1,000 images from the ImageNet validation set deng2012imagenet . This dataset has been frequently employed in related work dong2019evading ; gao2020patch ; xie2019improving ; yuan2022natural ; chen2024content .
Semantic Consistency Evaluation. We use CLIP Score radford2021learning , SSIM wang2004image , PSNR huynh2008scope , and LPIPS zhang2018unreasonable to measure the semantic consistency between the generated unrestricted adversarial examples and the clean image from multiple aspects.
Attack Evaluation. We compare SCA with other unrestricted adversarial attacks, including SAE hosseini2018semantic , ADef alaifari2018adef , cAdv bhattad2019unrestricted , tAdv bhattad2019unrestricted , ACE zhao2020adversarial , ColorFool shamsabadi2020colorfool , NCF yuan2022natural , AdvST 10292904 , and ACA chen2024content . The default parameter settings for these attacks are used. The evaluation metric employed is attack success rate (ASR), which measures the proportion of images misclassified by the target models.
@@
Implementation Details. All experiments are performed using PyTorch on an NVIDIA Tesla A100. The DDPM steps are set to $T=10\to 20$, with attack iterations $N_{a}=10$, $\eta=0.04$, $\kappa=0.1$, and $\mu=1$. The version of Stable Diffusion rombach2022high used is v1.5, and image captions are generated automatically via LLaVA-NeXT.
@@ 5.2 Semantic Consistency Comparison
0ptcleanRefer to caption 0ptcAdvRefer to caption
@@
0ptSAERefer to caption 0ptADefRefer to caption 0ptColorFoolRefer to caption
@@
0ptcleanRefer to caption 0ptcAdvRefer to caption 0ptSAERefer to caption 0ptADefRefer to caption 0ptColorFoolRefer to caption 0ptNCFRefer to caption 0ptAdvSTRefer to caption 0ptACARefer to caption 0ptOurs
@@
5.3 Comparison of Attack Success Rates
We systematically evaluate the adversarial robustness of convolutional neural networks (CNNs) and vision transformers (ViTs) using several prominent adversarial methods: SAE hosseini2018semantic , ADef alaifari2018adef , cAdv and tAdv bhattad2019unrestricted , ACE zhao2020adversarial , ColorFool shamsabadi2020colorfool , NCF yuan2022natural , AdvST 10292904 , ACA chen2024content , and our proposed SCA. These methods were applied to MobileNetV2 (MN-v2), ResNet-50 (RN-50), MobViT-s, and ViT-B, with performance measured using the Average Attack Success Rate (%), which captures the transferability of adversarial examples across non-substitute models.
@@ 5.3 Comparison of Attack Success Rates
5.4 Image Quality Comparison.
In this section, we conduct a detailed quantitative assessment of the image quality for adversarial examples, following the approach taken in prior works such as shamsabadi2020colorfool , yuan2022natural , and chen2024content . Specifically, we employ non-reference perceptual image quality metrics to evaluate the visual realism and aesthetic appeal of the adversarial images generated by different methods. The metrics we use include NIMA talebi2018nima , HyperIQA su2020blindly , MUSIQ ke2021musiq , and TReS golestaneh2022no , which are widely recognized for their ability to assess image quality in the absence of a ground truth reference.
To provide a robust evaluation, we follow the PyIQA framework and apply models trained on large-scale aesthetic datasets. NIMA-AVA and MUSIQ-AVA are trained on the AVA murray2012ava dataset, while MUSIQ-KonIQ is trained on the KonIQ-10K hosu2020koniq dataset, which ensures a comprehensive evaluation of the generated adversarial examples. As shown in Table 3, our method consistently achieves image quality scores that are comparable to ACA, outperforming other methods such as ColorFool and NCF. This superior performance is likely attributable to the fact that both SCA and ACA leverage Stable Diffusion, a state-of-the-art generative model known for its ability to produce highly realistic and visually appealing images.
@@
Additionally, it is worth noting that the no-reference image quality metrics we employ are often trained on aesthetic datasets, many of which contain post-processed images, including those edited with tools such as Photoshop. These post-processed images are more aligned with human aesthetic preferences, which is an important consideration for assessing adversarial examples. As chen2024content points out, the similarity of our method to post-processing and image-editing techniques helps explain why our adversarial examples receive higher quality scores from these metrics. Our approach effectively functions as a form of image editing, subtly modifying the image in ways that retain or even enhance its aesthetic appeal, leading to better evaluation results.